# PAMS: The Perseus Arm Molecular Survey – I. Survey description and first results

### v1

@@ 2.3 Ancillary data
We incorporated additional 12CO (=3–2) data covering from the JCMT archives covering seven out of the eight regions. G135.2+1.1, IRAS02327, S157, NGC 7538, and W5 have almost complete coverage, while G110 and S152 have partial coverage, and G134.9+1.4 has no coverage at all. The archival data used were the PI programs of Lumsden (M07AU08), Williams (M07BH45B, M08BH15, and M09BC12), Hogerheijde (M09BN07), Di Francesco (M09BH09C) and Reid (M10BC04). The NGC 7538 and W5 data were previously published in Fallscheer et al. (2013) and Ginsburg et al. (2011), respectively, while outflows within G110, S152, and S157 formed part of the sample of Maud et al. (2015).
These data were reduced in an almost identical way using orac-dr, differing only in the use of a 9-arcsec FWHM Gaussian smooth during the regridding, which results in data cubes whose effective angular resolution matches the 17.2-arcsecond resolution of the 13CO and C18O data cubes. Some of the CO data were observed with ACSIS configured with a 1000 MHz bandwidth rather than 250 MHz, which provides a native resolution of 0.42 km s-1, and so the CO data were universally rebinned to 0.5 km s-1 velocity channels throughout for consistency. These data were converted to the scale by dividing by the 345 GHz main beam efficiency of = 0.61 (Buckle et al., 2009). In Fig. 2 we displayed a combined view111Colour image created using the multicolorfits Python package. of the 12CO (3–2) emission alongside the PAMS 13CO (3–2) emission for NGC 7538, highlighting the utility of the combined data sets. Regions of orange emission trace the highest column densities of CO, while the yellow emission traces the diffuse envelope of the region.
@@ 3.1 Data quality
Fig. 3(a) shows the distributions of pixel values for the entirety of the PAMS 13CO and C18O (3–2) data, and in both the 0.3- and 0.5-km s-1 velocity channel-width variants. To determine a global rms value for each isotopologue, we fitted a normal distribution centred on a value of zero, and recovered rms values of 0.69 (0.54) and 0.89 (0.73) K in for 13CO and C18O at 0.3 (0.5) km s-1-resolution, respectively. Assuming abundance ratios of 12CO/13CO of 70 for sources at kpc (Milam et al., 2005) and 12CO/H2 of (Frerking et al., 1982), the rms in the 13CO data at 0.3 km s-1 resolution corresponds to a 13CO column density of cm-2 for optically thin gas at 10 K – equivalent to a H2 column density of cm-2. In all cases, the data are not perfectly normally distributed; the global rms values arise from the central limit theorem when combining a different distribution for each tile within each region. We display the values for each individual region in Table 1. In the inset axes of Fig. 3(a), the distributions are displayed with a logarithmic y-axis, and the non-Gaussian wings in the distributions are visible most clearly. The logarithmic 13CO distributions show a significant wing at positive values, which is associated with the emission. This is less obvious in C18O due to the lower relative abundance of the isotopologue compared to 13CO, which results in a much lower detection rate. The noise levels vary from tile to tile as result of different observing conditions: zenith opacity (i.e. precipitable water vapour), target elevation, and the number of functioning receptors on HARP. The tile-to-tile noise variations can be seen in Fig. 4, and we display the distributions of noise values for the various PAMS data sets in Fig. 3(b).
@@ 3.2 Emission maps
In Fig. 5, we present the observations of NGC 7538. We show both the moment 0 (velocity-integrated intensity) maps generated, and position-velocity maps (declination-integrated intensity) for each isotopologue of 12CO. The images were first masked using the FellWalker source extraction software (discussed in Section 3.3). Although the bulk of the emission from NGC 7538 is contained within 70 to 40 km s-1, the region contains a number of outflows that extend beyond this range. The outflows are clearly visible in the position-velocity maps of 12CO and, to a lesser extent, 13CO. The field also contains a secondary minor emission component between 17 and 3 km s-1.
With the lowest effective critical density, the 12CO (3–2) emission traces the most diffuse component of the molecular cloud, which fills much of the field of view, and it is especially powerful for tracing outflows. By contrast, the 13CO and C18O (3–2) emission traces higher-density components; 12CO is 80 and 600 times more abundant than 13CO and C18O, respectively, at a Galactocentric distance of 9.5 kpc (Wilson & Rood, 1994), and so the rarer isotopologues have substantially lower optical depths. Rigby et al. (2019) found that 13CO (3–2) emission within molecular clouds in the inner Milky Way is predominantly , and only becomes optically thick towards the densest 1 pc-scale clumps; only 3% of clumps in the CHIMPS sample have mean optical depths of greater than 1. NGC 7538 and the other PAMS regions also reside substantially further out in the Galaxy than the clumps typically targeted by CHIMPS, and so optical depths are likely to be even lower for PAMS. 13CO (3–2) emission is, therefore, expected to be a reasonably good tracer of H2 column density where it is detected in these regions.
In Fig. 5 we also overlay the loci of the models of the four spiral arms of Reid et al. (2019) the are present in this quadrant of the Galaxy. We clearly detect emission associated with the Local Arm in addition to the expected Perseus Arm – with which the main PAMS regions are associated. The Reid et al. (2019) models are based upon trigonometric parallaxes of 200 high-mass star-forming regions, obtained by the Bar and Spiral Structure Legacy (BeSSeL) Survey222http://bessel.vlbi-astrometry.org/ and the Japanese VLBI Exploration of Radio Astrometry (VERA) project333https://www.miz.nao.ac.jp/veraserver/ which are known to a high level of accuracy. NGC 7538 clearly resides within the Perseus arm, and the secondary emission component is consistent with a position in the Local Arm.
@@ 3.3 Source extraction
A series of source extractions was performed upon the 13CO (3–2) PAMS data using the FellWalker (Berry, 2015) algorithm, which is part of cupid (Berry et al., 2007). FellWalker is a watershed clump-finding algorithm which segments our arrays of three-dimensional pixels (or voxels) – with two spatial and one spectral dimension – into discrete clumps of emission which each contain a significant emission peak. FellWalker assigns all voxels brighter than a threshold level, determined by the noise, to a single clump in this way. In addition to producing a catalogue, it also generates a mask which has the same dimensions as the input array, but in which each pixel value corresponds to the ID of a catalogued clump. We refer to these masks as ‘assignment masks’ hereafter, and these may also be used to effectively remove the background from the data cubes, amongst other useful applications.
The source extraction was performed upon the signal-to-noise ratio (SNR) mosaics of each region. Initial tests showed that FellWalker was much more effective at locating the emission present within the data when running over the SNR cubes as opposed to the cubes, resulting in a smaller amount of emission remaining in the residual cubes. After an initial source-finding on the SNR cubes using cupid:findclumps, the cupid:extractclumps algorithm then uses the FellWalker-defined mask to extract information from the cubes, recalculating properties such as the intensity-weighted centroid and peak coordinate of each source, which is different in -space compared to SNR-space. A further advantage is that an extraction of sources on the SNR cubes reduces the instances of false positive detections which can arise as a consequence of noisy spectra at the image edges or at the seams of the mosaics.
@@
For both fwhires and fwlores, source extraction was performed on the 0.3-km s-1-resolution mosaics, essentially giving the ‘best’ case for our PAMS data, while the fwchimps extraction was conducted on the 0.5 km s-1-resolution mosaics to facilitate a more direct comparison with CHIMPS. These source extractions were also performed after smoothing the data to a resolution of 22.0 arcsec, which was found to be the best compromise between retaining relatively high resolution, while improving the noise statistics. For these extractions, the minimum height of a peak to be included was set to SNR = 3, and the noise level was set to SNR = 1, meaning that all contiguous pixels down to the rms value are considered to be part of each source. The difference in the fwhires and fwlores was achieved by setting different values of MinDip which was set to an SNR value of 5 in the fwhires case, and 1000 in the fwlores case. By contrast, the fwchimps extraction was configured to use the parameters described in (Rigby et al., 2016), with the exception that the RMS parameter was set to 1.7, reflecting the lower sensitivity of the CHIMPS data after smoothing to the same resolution of 27.4 arcsec that was originally used. One further difference in the process is that this extraction identifies sources at a resolution of 27.4 arcseconds, but then extracts the parameters at the native resolution of 17.2 arcsec (15.2 arcsec in the original CHIMPS extraction). Both PAMS extractions identify and extract source parameters at the same resolution of 22.0 arcsec. We list the full set of FellWalker parameters used for each setup in Appendix B.
@@
While FellWalker does not nominally contain any information about the hierarchical structure of the emission, our twin fwhires and fwlores allow some aspects of this to be recovered. Both extractions identify the same pixels of emission, and differ only in the assignment to catalogued structures. Because of this, we are able to assign every clump within the fwhires extraction to a larger structure from the fwlores extraction, and thus restore some information about the hierarchy. By contrast, the scimes (Colombo et al., 2015) algorithm is based upon the astrodendro implementation of the dendrograms (Rosolowsky et al., 2008), which identify substructures that are significant in terms of brightness and area based upon contour levels. While the dendrograms also keep track of compact sources (identified as leaves by analogy) inside larger substructures (branches) of molecular clouds (trunks), their individual values may differ compared to the equivalent FellWalker-defined clumps, but statistically the two approaches return compatible results (Rani et al., 2023). In this paper, we use FellWalker to enable a direct comparison with the properties of the clump population identified within the CHIMPS survey of Rigby et al. (2019) in the Inner Galaxy.
@@
Table 2 contains the first five rows of the fwhires catalogue, with selected information given. We make full versions of the fwlores, fwhires, and fwchimps catalogues available alongside this paper, and we detail the column descriptions in Appendix LABEL:app:table. The format of the three catalogues is almost identical, with the exception that the fwhires catalogue also lists the ID of the parent source in the fwlores catalogue to allow the hierarchical information about the larger complexes to be retained.
Fig. 6 illustrates the differences between the source extraction setups. The 2-D representations of the FellWalker masks clearly show the difference between the fwhires and fwlores extractions, illustrating that the same pixels of emission are recovered, but differ in their assignment to different structures. The fwchimps extraction shows the substantial difference that the data quality makes in source extraction, and the necessity to have a like-for-like extraction in order to make meaningful comparisons to other datasets that take into account biases resulting from sensitivity. Both fwhires and fwlores recover much fainter emission than fwchimps, which exists in the diffuse envelopes of the molecular clouds. The residual images give an idea of how complete the various extractions are. While fwhires and fwlores leave almost no visible residual in the integrated position-position intensity maps, those extractions still leave faint and unrecovered emission that is most clearly visible in the residual integrated position-velocity intensity maps.
@@ 4.1 CO-to-H2 conversion factors
The factor converts the integrated intensity of CO emission in a particular transition (and a particular isotopologue) to total molecular hydrogen column density. In its general form:
@@
where is the molecular hydrogen column density, and is the integrated intensity of 12CO (1–0) emission. For reference, Bolatto et al. (2013) recommend a typical value of cm-2 (km s-1)-1 in the Milky Way disc. condenses the wide range of environmentally-varying excitation conditions that are likely to be found within a relatively large area of Galactic disc (typically used in external galaxies) into a single scaling relationship. Here, we explore the value of for our 12CO and 13CO (3–2) emission lines (which we refer to as and , respectively) within PAMS, and compare the results to those derived from similar data in an Inner Galaxy field. The Inner Galaxy region we used is centred at °– which contains the massive star-forming complex W43 and cloud G29.960.02 – using mosaics of the 12CO COHRS (Dempsey et al., 2013; Park et al., 2023) and 13CO CHIMPS (3–2) data (Rigby et al., 2016), as re-processed for the CHIMPS2 Inner Galaxy survey (Rigby et al. in preparation) on 6 arcsec pixels, and with an effective resolution matching PAMS.
To calculate values, we compared the moment 0 (velocity-integrated intensity) maps to maps of H2 column density derived from the point process mapping (PPMAP) analysis of Herschel/Hi-GAL tiles from Marsh et al. (2017). The PPMAP technique combines surface brightness maps in the five Hi-GAL wavebands (70–500 m) with knowledge of the point spread functions for each waveband to produce maps of differential column density – evaluated at different dust temperatures – with an enhanced resolution of 12 arcseconds compared to more widely-used pixel-by-pixel greybody fitting. We make use of the PPMAP tiles of total column density that cover our regions, and also use the column density-weighted dust temperature maps for further analysis. Owing to the difference in latitude coverage of Hi-GAL compared to the PAMS regions, we could calculate factors only for G135, IRAS02327, NGC 7538, and the western section of W5 (we found a discrepancy with the astrometry in the PPMAP tile covering the eastern section).
@@
The 12CO and 13CO cubes, and H2 column density maps were first smoothed to a common angular resolution of 20 arcseconds, and the cubes were integrated over velocity ranges containing the emission to produce the moment 0 maps. We masked the moment 0 maps below a contour level determined for each region in order to limit the impact of noise. The H2 column density maps were then resampled onto the same pixel grid as the moment 0 images so that the column densities and integrated intensities could be compared on a pixel-by-pixel basis.
@@
We used Eq. 1, and performed a least-squares fit to the data points;
@@
We adapted the first approach to include a column density offset (i.e. background correction):
@@
Fig. 7 illustrates several aspects of these results. In panels a) and d), we show the distributions of pixel values of H2 column density as functions of the integrated intensity of 12CO and 13CO (3–2) in the corresponding pixel. In both cases, it is clear that the distributions are not completely linear (i.e. single power-laws in log-space). We see a flattening of the distributions that is more prominent in 12CO, and indicative of the column density background that is detected in the spaced-based Herschel observations, but not discernible in the ground-based CO observations due to their sensitivity. At the bright end of the distributions, the column density also curves upwards – and this behaviour is stronger in 12CO (3–2) – indicating a saturation of CO emission as a consequence of the emission becoming optically thick at the highest column densities. The global values are plotted as black lines, and these are the values that are representative of the kind of values that are used in studies of extragalactic systems, where star forming complexes may be unresolved. The scatter around the global values is clearly not random, and we explore the origin by examining the distributions as functions of dust temperature and column density in panels b), c), e), and f).
@@
The distributions of individual values are very broad, spanning between one and two orders of magnitude in both isotopologues. For all cases, the distributions of dust temperature are also broad for a given value. One very noticeable trend is that the lower envelope of the dust temperature distribution is 5 K higher in the Inner Galaxy sample (contours). It is also of particular note that the peak of the distribution of values in the Inner Galaxy is a factor of 2–3 higher than the global value for the Outer Galaxy, suggesting that global values of are likely to be weighted towards particular ISM conditions that are conducive to bright CO emission, but which are not representative of typical conditions by area.
@@
For , the range of values returned from our different methods are fairly comparable, somewhere in the range of 2– cm-2 (km s-1)-1. Each of the methods systematically returns a higher aggregate value of for the Outer Galaxy than for the Inner Galaxy region. However, the difference is small, being marginally (15%) higher in the Outer Galaxy. Given the different calculation methods, we suggest that our method i) values are most appropriate for resolved studies, and method iv) values are more appropriate for unresolved studies. We therefore recommend the usage of a value of cm-2 (km s-1)-1 for resolved studies and cm-2 (km s-1)-1 for unresolved studies. In the case of 12CO (3–2), these values are almost identical, but differ somewhat from the value of cm-2 (km s-1)-1 derived by Colombo et al. (2019). The latter value was derived by applying an average 3–2/1–0 line ratio measurement to the Bolatto et al. (2013) value listed above and, given that the approach is somewhat different to ours, a discrepancy is not surprising. We benchmark our values using the PPMAP-derived Herschel column densities of Marsh et al. (2017) whose sophisticated approach (particularly allowing for multiple line-of-sight dust temperatures) ought to yield reasonably accurate results.
@@
The picture is more varied for , with values ranging from 1.3 to 4.5 cm-2 (km s-1)-1 between the different regions and calculation methods. Three of the four methods return higher values in the Inner Galaxy than the Outer Galaxy, with method ii) that includes a fitted background value being the exception. The higher column density background in the Inner Galaxy, which arises as a consequence of the greater number of spiral arms present along the line of sight and a greater column of Galactic disc, appears to be the primary driver of the different recovered values. This variation is clearly illustrated in Fig. 7d where the global average lines (for method iv)) are more widely separated for the Inner and Outer Galaxy. The peak of the distributions in Figs. 7e and 7f between the Inner and Outer Galaxy are noticeably different. In the Inner Galaxy, it appears that the higher dust-temperature background of 20 K compared to 15 K in the Outer Galaxy, along with a higher background of column density contribute to the higher values. The single dust temperature is less likely to represent the underlying gas conditions at this Inner Galaxy position, where a greater fraction of the Galactic disc is located within the column. We note that gas and dust temperatures are only likely to be coupled at densities greater than cm-3 (Goldsmith, 2001), so that for much of the gas in molecular clouds – especially their envelopes – the CO excitation temperature is unlikely to follow the dust temperature. We will explore CO excitation temperatures, along with a local thermodynamic equilibrium (LTE)-based derivation in a future paper.
@@
For resolved studies, we recommend the usage of our method i) value of cm-2 (km s-1)-1 as a representative value for , and for unresolved studies, we recommend a value of cm-2 (km s-1)-1. As for 13CO, these factors should always be used with the knowledge that their derivation is biased towards (relatively) hot high-column density gas where CO is brightest on large scales, but will not accurately account for column densities in other environments. In all cases, we recommend that a multiplicative factor of 1.5 be adopted for the uncertainty on the values to encapsulate the 1- spread of the values.
4.2 Molecular cloud properties

### v3

@@ 2.5 Ancillary data
We make use of the data from the CHIMPS survey (Rigby et al., 2016) directly in Section 4.1, and use the clump catalogue from Rigby et al. (2019) in Section 4.2.
We also use 160 and 250 m data from Hi-GAL (Molinari et al., 2016) to construct maps of H2 column density and dust temperature in Section 4.1. These data were reduced using UniMap – the University of Rome and IAPS Map Maker – version 5.1.0 (Molinari et al. in prep), which include de-striping, and the application of photometric offsets determined by comparison to Planck data.
@@ 3.1 Data quality
Fig. 3 shows the distributions of voxels (i.e. three-dimensional pixel) values for the entirety of the PAMS 12CO, 13CO, and C18O (3–2) data, and in both the 0.3- and 0.5-km s-1 velocity channel-width variants. To determine a global rms value for each isotopologue, we fitted a normal distribution centred on a value of zero, and recovered rms values of 1.00 (0.72), 0.68 (0.53), and 0.88 (0.71) K in for 12CO, 13CO, and C18O at 0.3 (0.5) km s-1-resolution, respectively. The RMS-equivalent H2 column densities are , , and cm-2 for 12CO, 13CO, and C18O (3–2), respectively, assuming optically thin emission with an excitation temperature of 10 K. For these figures, we have assumed abundance ratios of for 12CO/H2 (Frerking et al., 1982), and 79 and 596 for 12CO/13CO and 12CO/C18O, respectively, from Wilson & Rood (1994) evaluated at kpc. In all cases, the data are not perfectly normally distributed; the global rms values arise from the central limit theorem when combining a different distribution for each tile within each region. We display the values for each individual region in Table 1. In Fig. 3, the logarithmic -axes allow the non-Gaussian wings of the distributions to be seen most clearly, while the linear -axes in the inset figures demonstrate the overall normal distributions. In 12CO and 13CO, the distributions show significant positive wings associated with the emission. This is less obvious in C18O due to the lower relative abundance of the isotopologue, resulting in a much lower detection rate. The noise levels vary from tile to tile as a result of different observing conditions: zenith opacity (i.e. precipitable water vapour), target elevation, and the number of functioning receptors on HARP. The tile-to-tile noise variations can be seen in Fig. 4, and we display the distributions of noise values for the various PAMS data sets in Fig. 3.
@@ 3.2 Emission maps
In Fig. 5, we present the observations of NGC 7538, and equivalent figures for the other regions are shown in Appendix B. We show both the moment 0 (velocity-integrated intensity) maps, and position-velocity maps (declination-integrated intensity) generation for each isotopologue of 12CO. The images were first masked using the FellWalker source-extraction software (discussed in Section 3.3). Although the bulk of the emission from NGC 7538 is contained within 70 to 40 km s-1, the region contains a number of outflows that extend beyond this range. The outflows are clearly visible in the position-velocity maps of 12CO and, to a lesser extent, 13CO. The field also contains a minor secondary emission component between 17 and 3 km s-1, which contains just 8 and 1 per cent of the total emission in 12CO and 13CO (3–2), respectively.
With the lowest effective critical density, the 12CO (3–2) emission traces the most diffuse component of the molecular cloud, which fills much of the field of view, and it is especially powerful for tracing outflows. By contrast, the 13CO and C18O (3–2) emission traces higher-density components; 12CO is 80 and 600 times more abundant than 13CO and C18O, respectively, at a Galactocentric distance of 9.5 kpc (Wilson & Rood, 1994), and so the rarer isotopologues have substantially lower optical depths. Rigby et al. (2019) found that 13CO (3–2) emission within molecular clouds in the inner Milky Way is predominantly , and only becomes optically thick towards the densest 1 pc-scale clumps; only 3 per cent of clumps in the CHIMPS sample have mean optical depths greater than 1. NGC 7538 and the other PAMS regions also reside substantially further out in the Galaxy than the clumps typically targeted by CHIMPS, and so optical depths are likely to be even lower for PAMS. 13CO (3–2) emission is, therefore, expected to be a reasonably good tracer of H2 column density where it is detected in these regions.
In Fig. 5 we also overlay the loci of the models of the two spiral arms of Reid et al. (2019) that are present in this quadrant of the Galaxy, which fall within the PAMS latitude range. We clearly detect emission associated with the Local Arm in addition to the expected Perseus Arm – with which the main PAMS regions are associated. The Reid et al. (2019) models are based upon trigonometric parallaxes of 200 high-mass star-forming regions, obtained by the Bar and Spiral Structure Legacy (BeSSeL) Survey222http://bessel.vlbi-astrometry.org/ and the Japanese VLBI Exploration of Radio Astrometry (VERA) project333https://www.miz.nao.ac.jp/veraserver/ which are known to a high level of accuracy. NGC 7538 clearly resides within the Perseus arm, and the secondary emission component is consistent with a position in the Local Arm.
@@ 3.3 Source extraction
A series of source extractions was performed upon the 13CO (3–2) PAMS data using the FellWalker (Berry, 2015) algorithm, which is part of cupid (Berry et al., 2007). FellWalker is a watershed clump-finding algorithm which segments our arrays of voxels into discrete clumps of emission, each of which contains a significant emission peak. FellWalker assigns all voxels brighter than a threshold level, determined by the noise, to a single clump in this way. In addition to producing a catalogue, it also generates a mask which has the same dimensions as the input array, but in which each voxel value corresponds to the ID of a catalogued clump. We refer to these masks as ‘assignment masks’ hereafter, and these may also be used to effectively remove the background from the data cubes, amongst other useful applications.
Source extraction was performed upon the signal-to-noise ratio (SNR) mosaics of each region. Initial tests showed that FellWalker was much more effective at locating the emission present within the data when running over the SNR cubes as opposed to the cubes, resulting in a smaller amount of emission remaining in the residual cubes. After an initial source search on the SNR cubes using cupid:findclumps, the cupid:extractclumps algorithm then uses the FellWalker-defined mask to extract information from the cubes, recalculating properties such as the intensity-weighted centroid and peak coordinate of each source, which is different in -space compared with SNR-space. A further advantage is that an extraction of sources on the SNR cubes reduces the instances of false-positive detections which can arise as a consequence of noisy spectra at the image edges or at the seams of the mosaics.
@@
For both fwhires and fwlores, source extraction was performed on the 0.3-km s-1-resolution mosaics, essentially giving the ‘best’ case for our PAMS data, while the fwchimps extraction was conducted on the 0.5 km s-1-resolution mosaics to facilitate a more direct comparison with CHIMPS. These source extractions were also performed after smoothing the data to a resolution of 22.0 arcsec, which was found to be the best compromise between retaining relatively high resolution, while improving the noise statistics. For these extractions, the minimum height of a peak to be included was set to SNR = 3, and the noise level was set to SNR = 1, meaning that all contiguous pixels down to the rms value are considered to be part of each source. Although the noise level (i.e. minimum voxel value to be included in a clump) is rather low, we found that higher levels for this parameter led to considerably more emission in the residuals. As always, flux boosting will be present in the low-SNR sources, and we stress that this extraction has been tuned to maximise the recovery of the emission, as opposed to catalogue robustness (i.e. minimizing false positives). The difference in the fwhires and fwlores was achieved by setting different values of MinDip which was set to an SNR value of 5 in the fwhires case, and 1000 in the fwlores case. By contrast, the fwchimps extraction was configured to use the parameters described in (Rigby et al., 2016), with the exception that the RMS parameter was set to 1.7, reflecting the lower sensitivity of the CHIMPS data after smoothing to the same resolution of 27.4 arcsec that was originally used. One further difference in the process is that this extraction identifies sources at a resolution of 27.4 arcsec, but then extracts the parameters at the native resolution of 17.2 arcsec (15.2 arcsec in the original CHIMPS extraction). Both PAMS extractions identify and extract source parameters at the same resolution of 22.0 arcsec. We list the full set of FellWalker parameters used for each setup in Appendix C.
@@
While FellWalker does not nominally contain any information about the hierarchical structure of the emission, our twin fwhires and fwlores allow some aspects of this to be recovered. Both extractions identify the same pixels of emission, and differ only in the assignment to catalogued structures. Because of this, we are able to assign every clump within the fwhires extraction to a larger structure from the fwlores extraction, and thus restore some information about the hierarchy. By contrast, the scimes (Colombo et al., 2015) algorithm is based upon the astrodendro implementation of dendrograms (Rosolowsky et al., 2008), which identify substructures that are significant in terms of brightness and area based upon contour levels. While dendrograms also keep track of compact sources (identified as leaves by analogy) inside larger substructures (branches) of molecular clouds (trunks), their individual values may differ compared with the equivalent FellWalker-defined clumps, but statistically the two approaches return compatible results (Rani et al., 2023). In this paper, we use FellWalker to enable a direct comparison with the properties of the clump population identified within the CHIMPS survey of Rigby et al. (2019) in the Inner Galaxy.
@@
Table 2 contains the first five rows of the fwhires catalogue, with selected information given. We make full versions of the fwlores, fwhires, and fwchimps catalogues available alongside this paper, and we detail the column descriptions in Appendix D. The format of the three catalogues is almost identical, with the exception that the fwhires catalogue also lists the ID of the parent source in the fwlores catalogue to allow the hierarchical information about the larger complexes to be retained.
Fig. 6 illustrates the differences between the source- extraction setups. The two-dimensional representations of the FellWalker masks clearly show the difference between the fwhires and fwlores extractions, illustrating that the same pixels of emission are recovered, but differ in their assignment to different structures. The fwchimps extraction shows the substantial difference that the data quality makes in source extraction, and the necessity to have a like-for-like extraction in order to make meaningful comparisons to other data sets that take into account biases resulting from sensitivity. Both fwhires and fwlores recover much fainter emission than fwchimps, which exists in the diffuse envelopes of the molecular clouds. The residual images give an idea of how complete the various extractions are. While fwhires and fwlores leave almost no visible residual in the integrated position-position intensity maps, those extractions still leave faint and unrecovered emission that is most clearly visible in the residual integrated position-velocity intensity maps.
@@ 4.1 CO-to-H2 conversion factors
The factor converts the integrated intensity of CO emission in a particular transition (and a particular isotopologue) to total molecular-hydrogen column density. In its general form:
@@
where is the molecular-hydrogen column density, and is the integrated intensity of 12CO (1–0) emission. For reference, Bolatto et al. (2013) recommend a typical value of cm-2 (K km s-1)-1 in the Milky Way disc. condenses the wide range of environmentally varying excitation conditions that are likely to be found within a relatively large area of Galactic disc (typically used in external galaxies) into a single scaling relationship. Here, we explore the value of for our 12CO and 13CO (3–2) emission lines (which we refer to as and , respectively) within PAMS, and compare the results with those derived from similar data in an Inner Galaxy field. The Inner Galaxy region we used is centred at °– which contains the massive star-forming complex W43 and cloud G29.960.02 – using mosaics of the 12CO COHRS (Dempsey et al., 2013; Park et al., 2023) and 13CO (3–2) CHIMPS data (Rigby et al., 2016). The latter were re-processed for the CHIMPS2 Inner Galaxy survey (Rigby et al. in preparation) on 6 arcsec pixels, with an effective resolution matching PAMS.
To calculate values, we compared the moment 0 (velocity-integrated intensity) maps to maps of H2 column density derived from greybody fitting with 160 and 250 m imaging from the Herschel/Hi-GAL survey following the method of Peretto et al. (2016). The specific intensity at a frequency is related to H2 column density via
where is the mean molecular mass per H2 (equal to 2.8 for molecular gas with a relative Helium abundance of 25 per cent), is the mass of a hydrogen atom, is the gas-to-dust mass ratio, is the dust opacity at a reference frequency of , is the value of the Planck function evaluated at frequency for the dust temperature . First, maps of dust temperature are determined from the ratio of 160 to 250 m flux densities after convolving the former to 18-arcsec resolution to match the latter:
for which we adopt , and sample dust temperatures to the nearest 0.01 K. We then generate maps of column density using the corresponding 250 m map:
using a value of cm2 g-1 at 250 m444The dust opacity is interpolated from Table 1, column 5 of Ossenkopf & Henning (1994), but scaled down by a factor of 1.5, which was found by Kauffmann et al. (2010) to better reproduce extinction measurements., and a typical gas-to-dust mass ratio of .
@@
The 12CO and 13CO cubes, and H2 column-density maps were then smoothed to a common angular resolution of 20 arcsec, and the cubes were integrated over velocity ranges containing the emission to produce moment 0 (integrated intensity) maps. We masked the moment 0 maps below a contour level determined for each region in order to limit the impact of noise. The H2 column-density maps were then resampled onto the same pixel grid as the moment 0 images so that the column densities and integrated intensities could be compared on a pixel-by-pixel basis. The overlap between Hi-GAL and PAMS meant that we could calculate factors for around 70 per cent of the survey area with S152, S157, and most of G110 falling out of the Hi-GAL coverage (see Fig. 1).
@@
We used equation (1), and performed a least-squares fit to the data points;
@@
We adapted the first approach to include a column-density offset (i.e. background correction):
@@
We show the distributions of pixel values across PAMS in Fig. 7, in which it is apparent that the Outer Galaxy values are lower than the Inner Galaxy values for both 12CO and 13CO (3–2), with the discrepancy being larger for the latter. Fig. 8 illustrates several additional aspects of these results by exploring the correlation between individual pixel values and other properties. In Panels a) and d), we show the distributions of pixel values of H2 column density as functions of the integrated intensity of 12CO and 13CO (3–2) in the corresponding pixel. In both cases, it is clear that the distributions are not completely linear (i.e. simple power-laws in log-space) and there is considerable scatter. We see a flattening of the distributions at low integrated intensity that is more prominent in 12CO, and indicative of the column-density background that is detected in the spaced-based Herschel observations, but not discernible in the ground-based CO observations due to their sensitivity. At the bright end of the distributions, the column density also curves upwards – and this behaviour is stronger in 12CO (3–2) – indicating a saturation of CO emission as a consequence of the emission becoming optically thick at the highest column densities. The global values (Method iv) are plotted as black lines, and these are the values that are representative of the kind of values that are used in studies of extragalactic systems, where star-forming complexes may be unresolved. The scatter around the global values is clearly not random, and we explore the origin by examining the distributions as functions of dust temperature and column density in Panels b), c), e), and f).
@@
The distributions of individual values are very broad, spanning between one and two orders of magnitude in both isotopologues. For all cases, the distributions of dust temperature are also broad for a given value. One very noticeable trend is that the lower envelope of the dust temperature distribution is 5 K higher in the Inner Galaxy samples (contours). It is also of particular note that the peak of the distribution of values in the Inner Galaxy is a factor of 2–3 higher than the global value for the Outer Galaxy, suggesting that global values of are likely to be weighted towards particular ISM conditions that are conducive to bright CO emission, but which are not representative of typical conditions by area.
@@
For , the range of values returned from our different methods are fairly comparable, ranging between 2.3–4.2 cm-2 (K km s-1)-1. Each of the methods systematically returns a higher value of for the Inner Galaxy than for the Outer Galaxy region although the distributions overlap substantially (see Fig. 7). Given the different calculation methods, we suggest that – in principle – our Method i) values are most appropriate for resolved studies, and Method iv) values are more appropriate for unresolved studies. Given the insignificant difference for 12CO (3–2), we recommend the usage of = cm-2 (K km s-1)-1 for both resolved studies and unresolved studies, noting that range of values varies by around per cent (as per Method iii). This is in excellent agreement with Colombo et al. (2019) who derived by their value by applying an average 3–2/1–0 line-ratio measurement to the Bolatto et al. (2013) value listed above using the wider COHRS data set of which we take a subset for our Inner Galaxy sample.
@@
The picture is more varied for , with values between 2.2–7.0 cm-2 (K km s-1)-1 across the different regions and calculation methods. All four methods return higher values in the Inner Galaxy than in the Outer Galaxy by a factor of 2–3. The higher column-density background in the Inner Galaxy, which arises as a consequence of the greater number of spiral arms present along the line of sight and a greater column of Galactic disc, appears to be the primary driver of the different recovered values. This variation is clearly illustrated in Fig. 8d where the global average lines (for Method iv) are more widely separated for the Inner and Outer Galaxy. The peak of the distributions in Figs. 7, 8e, and 8f between the Inner and Outer Galaxy are noticeably different. In the Inner Galaxy, it appears that the higher dust-temperature background of 20 K compared with 17 K in the Outer Galaxy, along with a higher background of column density contribute to the higher values. The single dust temperature is less likely to represent the underlying gas conditions at this Inner Galaxy position, where a greater fraction of the Galactic disc is located within the column. We note that gas and dust temperatures are only likely to be coupled at densities greater than cm-3 (Goldsmith, 2001), so that for much of the gas in molecular clouds – especially their envelopes – the CO excitation temperature is unlikely to follow the dust temperature. We will explore CO excitation temperatures, along with a local thermodynamic equilibrium (LTE)-based derivation in a future paper.
@@
For resolved studies, we recommend the usage of our Method i) value of cm-2 (K km s-1)-1 as a representative value for , and for unresolved studies, we recommend a value of cm-2 (K km s-1)-1. As for 12CO, these factors should always be used with the knowledge that their derivation is biased towards (relatively) hot high-column-density gas where CO is brightest on large scales, but will not accurately account for column densities in other environments. It is also important to recognise that a Galactic gradient of values is evident in our data, which we will discuss further in Section 5. In all cases, we recommend that a multiplicative factor of 1.5 be adopted for the uncertainty in the values to encapsulate the 1- spread of the values.
4.2 Molecular-cloud properties